OpenStudy (anonymous):

Beverly can drive 600 miles in the same time as it takes susan to drive 500 miles. If beverly drives 10 mph faster than susan, then how fast does beverly drive? What would the algebraic equation setup for this problem be?

OpenStudy (radar):

Let x = rate of Susan in mph, then x+10 is the rate of Beverly. d=rt distance equals rate times time. t=d/r 600/(x+10) =500/x t=t

OpenStudy (anonymous):

Thank you very much it's all clear now.

OpenStudy (radar):

Solving by cross multiplying getting 600x=500x+5000 100x=5000 x=50 mph x+10 =60 mph

OpenStudy (radar):

ur welcome,